OpenStudy (anonymous):

Matrices help? Let A = [1] and B = [-1] find A-3B. [-3] [3] [2] [-2]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

(1)x(-3)= (-3)x(-1)=3

OpenStudy (anonymous):

A and B are both 3x1 matrices. My answers are also 3x1 matices.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

I'm not sure how else I can show them. But in A 1 is the top number, -3 is the middle number, and 2 is the bottom number. In B -1 is the top number, 3 is the middle number, and -2 is the bottom number.

OpenStudy (anonymous):

Its a little hard to understand, sorry i don't know.

OpenStudy (anonymous):

for the matrix first, you can say ordo of the matrix, then for the matrix elements you can read it by row

OpenStudy (anonymous):

\[A = \left(\begin{matrix}1 \\ \begin{matrix}-3 \\ 2\end{matrix}\end{matrix}\right) ???\]

OpenStudy (anonymous):

\[B = \left(\begin{matrix}-1 \\ \begin{matrix}3 \\ -2\end{matrix}\end{matrix}\right) ???\]

OpenStudy (anonymous):

whether the matrix you mean like that?

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ \begin{matrix}-3 \\ 2\end{matrix}\end{matrix}\right) - 3 \left(\begin{matrix}-1 \\ \begin{matrix}3 \\ -2\end{matrix}\end{matrix}\right)\]

OpenStudy (anonymous):

to 3B, means multiply 3 to all elements of the matrix B